# LatticeVision: Image to Image Networks for Modeling Non-Stationary Spatial Data

### v1

@@ 4 Parameter estimation networks
We adapt three I2I networks for parameter estimation: a fully convolutional U-Net [ronneberger2015u], a modified ViT [dosovitskiy2020image], and a hybrid network inspired by the TransUNet [chen2021transunet] architecture. In each case we adhere closely to the original design, making only minimal or task-specific changes. A range of CNNs representative of the local estimation literature serve as our baselines.
UNet.
We use a standard U-Net [ronneberger2015u] with a symmetric encoder-bottleneck-decoder structure and skip connections that propagate spatial information across resolutions. Our implementation replaces ReLU with GELU activations [gelu], employs group normalization [groupnorm], and omits dropout [srivastava2014dropout]. The number of channels in the bottleneck matches the transformer embedding dimension used in the transformer-based I2I networks.
ViT.
The original ViT was designed for image classification [dosovitskiy2020image], so we remove classification-specific components such as the class token and final MLP head, and use a learnable linear layer to project the transformer output back to the original image resolution. The original 1D positional embeddings are replaced with a range of 2D alternatives; see Section 5.
STUN.
Our hybrid architecture, a spatial TransUNet (STUN), is based on the network in [chen2021transunet], which was originally designed for image segmentation. It blends the two I2I networks discussed above, originally combining a pretrained convolutional encoder with a vision transformer and a shallow decoder. We retain the overall structure, but replace the encoder and decoder with the symmetric U-Net components used in our fully convolutional network.
CNN.
Our local estimators follow the standard design of convolutional layers and a max pooling layer followed by an MLP. We use three networks with varying receptive window sizes that are representative of the local estimation literature [banesh2021fast, lenzi2023neural, gerber2021fast, sainsbury2024likelihood, wiens2020modeling]. The key difference is our networks have more parameters (1-2.5M), as compared to typical local estimators (50-700k).
@@ 4.1 Implementation and training details
All networks are trained with a batch size of $b=64$ and varying numbers of replicate input fields $M\in\{1,5,15,30\}$. Unless otherwise noted, all reported network sizes and metrics correspond to $M=30$ replicates. For the I2I networks, each training example consists of $M$ input fields of shape $[b,M,H,W]=[64,M,192,288]$, with corresponding output parameter fields $[64,3,192,288]$. For the local CNN estimators, we vary the receptive field $h=w=\{9,17,25\}$, with inputs of shape $[64,M,h,w]$ and scalar outputs $[64,3]$. We append the window size to the name of the CNN, thus $h=w=25$ is denoted CNN25. To encourage permutation invariance, we randomly shuffle the replicates across the channel dimension each training step, imposing a soft constraint. This avoids sacrificing network expressiveness, which is a concern when imposing a hard constraint via an aggregation step [deepsets]. We take care to preserve key elements of the statistical structure of the images, which are not compatible with data augmentation techniques such as image rotation, which does not preserve the angle of anisotropy $\theta$. Thus, we limit augmentation to spatial translation and random field negation. We generate datasets consisting of 8000 (I2I) and 80000 (CNN) samples with 30 replicates, and employ a 90/8/2 (train/validation/test) split, resulting in a test set of 160 samples for the I2I dataset, which both local and global methods are evaluated on in the following section. More details on implementation, storage, and software can be found in Appendix C.
@@ 5 Simulated data experiments
We generate a test set of $160$ samples following the procedure in Section 3. Each sample is a small ensemble $Y\in\mathbb{R}^{M\times 192\times 288}$ of $M$ replicates and an associated parameter image $\Phi\in\mathbb{R}^{3\times 192\times 288}$. For the I2I networks (UNet, ViT, STUN) we simply forward–propagate $Y$ to obtain $\hat{\Phi}_{\text{I2I}}$. For the local CNN baselines, we employ a pixel-by-pixel approach: we translate the CNN window across the field with a stride of 1, assigning the prediction to the central pixel of the window to build $\hat{\Phi}_{\text{CNN}}$. We use reflection padding---improving upon prior works which use zero padding---thus reducing edge-artifacts. Figure 4 contrasts an example $\Phi$ with estimates from the best performing global (STUN) and local (CNN25) networks. MLE is not attempted; local likelihood evaluation for one field would require hundreds of hours, exceeding the total runtime of our approach by orders of magnitude [wiens2020modeling].
Figure 4: True parameters $\Phi$ (left) are encoded into the LatticeKrig SAR model to simulate a testing sample $Y$, of which one replicate $\mathbf{y}^{(0)}$ is displayed. $Y$ is used as an input to STUN and a sliding window, local estimation strategy using CNN25, resulting in $\hat{\Phi}_{\text{STUN}}$ (middle), and $\hat{\Phi}_{\text{CNN25}}$ (right).
Positional embeddings.
The spatial nature of our problem leads us to study the effects of positional embeddings. We evaluate four embeddings for the transformer: none, 2-D sinusoidal [parmar2018image], learned 2-D [dosovitskiy2020image], and rotary (RoPE) [su2024roformer]. RoPE yields the best performance across the majority of metrics for both ViT and STUN, and is adopted henceforth (see Appendix Table 2).
Number of replicates.
@@
Table 1: Root mean square error (RMSE) for parameter estimation on simulated test set using 1 and 30 replicates (reps). ‘‘Size’’ is the net parameter count (in millions). ‘‘Train’’ is the wall-clock time for training until early stopping, and ‘‘Eval’’ is the average time (out of 5) to process the $160$ sample test dataset. Arrows indicate desirable direction, and bold values indicate best performance.
Net 30 rep RMSE $\downarrow$ 1 rep RMSE $\downarrow$ Size (M)$\downarrow$ Train (min)$\downarrow$ Eval (sec)$\downarrow$
@@
Our experiments highlight three consistent trends. (i) Global I2I networks pose a significant improvement over local CNNs: STUN and UNet often display 4-5x lower RMSE and are almost unaffected by shrinking the ensemble size, whereas CNNs are noticeably sensitive. (ii) Adding attention helps only marginally in this setting, and attention on its own lags behind approaches that include multiscale convolutions. STUN edges out UNet but at 4x the parameter count. ViT lags behind the other I2I networks, perhaps as it requires more training data [dosovitskiy2020image]. (iii) Global I2I networks take longer to train due to a higher parameter count, yet are amortized at inference time. In order to achieve the results of a single forward pass through an I2I network, the local estimators must perform $H\times W=55,296$ forward passes. Consequently, I2I networks perform inference 100--1000 times faster than local estimators.
@@ 6 Climate application
We evaluate our framework on a 30-member ensemble of surface temperature sensitivity fields from the Community Earth System Model Large Ensemble (CESM1 LENS) project [kay2015community]. Each field is defined on a grid of approximately one degree resolution ($192\times 288$), and represents the local change in temperature given an increase of $1^{\circ}$ in global temperature. Due to the chaotic nature of the ocean-atmosphere system, we can treat these 30 estimates of temperature sensitivity as independent replicates from the ‘‘true’’ climate sensitivity of CESM1. The data is preprocessed into standardized temperature sensitivity anomalies where each pixel has zero mean and unit variance. We then perform parameter estimation with the best global (STUN) and local (CNN25) estimation networks. Regardless of the estimation method used, our SAR model simulates $1000$ fields in less than $10$ seconds on a single laptop CPU, a stark contrast to the tens of millions of core hours required to generate the original 30-member ensemble.
@@
Figure 5: Top: Standardized temperature fields drawn from the CESM1 LENS ensemble (left), the STUN-based emulator (middle), and the CNN25-based emulator (right). Bottom: Correlations with a chosen location in the Niño 3.4 region at ($212^{\circ}$, $1^{\circ}$) for the same three ensembles. The STUN-based emulator better preserves spatial relationships, including the zonal correlation structure along the equator and the meridional oceanic correlation range.
@@
We use the STUN-based and CNN25-based emulators to generate 1000 fields each, and evaluate how well these synthetic ensembles preserve the spatial relationships present in the CESM outputs. Absolute prediction error is not meaningful here as the ‘‘truth’’ is itself a Monte-Carlo sample of internal climate variability. Thus, we compare the second-order structure through a covariance analysis. Rather than computing the entire $296\times 55,$ 296 × 55 , 296 correlation matrices of the true and simulated fields, we compare a representative sample of $50$ rows. First, we randomly select 50 anchor locations. For each of the three ensembles, we then calculate each anchor’s Pearson correlation with every other location. The average RMSE between these 50 true rows and their simulated counterparts is $0.229$ for the STUN-based and $0.243$ for CNN25-based emulation. A paired t-test demonstrates the strong statistical significance ($p_{\text{value}}=1.1\times 10^{-6}<0.01=\alpha$) of this difference (full results in Appendix E). Visualizing the empirical correlations with a chosen point in the Niño3.4 region---a region of the tropical pacific most correlated with the El Niño-Southern Oscillation phenomenon [barnston1997documentation]---provides a qualitative comparison of the two methods (Figure 5). The STUN-based emulator better preserves the expected zonal east-west correlation structure along the equator with a more realistic correlation range. The CNN-based emulator systematically over-smooths and inflates the oceanic correlation range, particularly in the meridional north-south direction. The original ensemble correlation field is noisier due to the CESM1 LENS ensemble only containing 30 members. In sum, the STUN method outperforms the CNN method both quantitatively and qualitatively in representing the underlying correlation structure of climate sensitivity fields.
@@ 7 Related works
Neural parameter estimation.
In situations where the likelihood function is intractable, but simulation from the model is feasible, neural networks have emerged as a powerful alternative [liu2020task, zammit2024neural] to MLE [wiens2020modeling]. Neural parameter estimation combines ideas from simulation-based inference [cranmer2020frontier, sisson2018handbook] and learning to optimize (L2O) [chen2022learning, yin2022learning] by training networks to identify maxima of the Bayes risk [sainsbury2024likelihood]. The training cost is then amortized by repeated use. Until now, these approaches have been limited to local estimation [banesh2021fast, lenzi2023neural, sainsbury2024likelihood, rai2024fast, gerber2021fast, rai2025extmodeling], typically using multi-layer perceptrons (MLPs) and CNNs. This line of research is the main inspiration for this work, which advances the field with simultaneous, non-stationary parameter inference and the use of I2I networks. Our network choices [ronneberger2015u, dosovitskiy2020image, chen2021transunet] establish a baseline, and there exist many additional architectures for future work to explore [cao2022swin, rao2022hornet, kim2022instaformer, wang2021pyramid, bao2023channel].
Climate model emulation.
While climate model emulation is not the sole application of our method, it provides an illustrative example and grounds our discussion of related, non-stationary emulation literature. Physics-based climate models require tens to hundreds of millions of core hours, necessitating data-driven emulation [kashinath2021physics, eyring2024pushing]. Despite recent advances in deterministic [pathak2022fourcastnet, lam2023learning, nguyen2024scaling, bi2023accurate, nathaniel2024chaosbench, keisler2022forecasting, chen2023fengwu] and ensemble-based probabilistic [kochkov2024neural, price2023gencast, li2023seeds, shi2024codicast] weather forecasting, comparatively fewer methods focus explicitly on long-term climate projections [lai2024machine, kashinath2021physics, eyring2024pushing]. Our approach draws ideas from statistical methods, which explicitly link parameter estimation with emulation [castruccio2014statistical, song2024efficient, nychka2018modeling, wiens2020modeling, chakraborty2025learning], and machine learning (ML) methods, both deterministic [nguyen2023climax, watt2023ace, chapman2025camulator] and probabilistic [ruhling2024probablistic, ruhling2023dyffusion] which emulate directly from initial forcings or prior timesteps. Specifically, we combine the straightforward uncertainty quantification and interpretable parameters from the statistical model with a deterministic ML approach for efficient parameter estimation. Purely ML-based climate emulators typically require an extensive corpus of training data [watson2022climatebench, yu2023climsim] and can exhibit limited generalization capabilities beyond their training distributions [kaltenborn2023climateset]. We sidestep this by generating synthetic training data that mimics non-stationarity in geophysical settings. Our framework does not serve as a replacement for ESMs, but rather a complementary method for augmenting ensemble sizes, and reducing the number of ‘‘ground truth’’ ESM runs that must be computed.
@@ 8 Limitations and extensions
Estimation networks.
Our I2I networks assume complete, regularly gridded data and are trained on fixed input dimensions. While the architectures can process variable resolutions, generalization across sizes remains untested. Replicate count is also fixed at training time, with only a soft constraint to enforce permutation invariance. A hard constraint via aggregation [deepsets] could resolve this, although the potential to limit network expressiveness must be explored. The performance of the transformer-based networks would likely benefit from efficient training regimes [touvron2021training] and a larger dataset. Future work could extend the data generation pipeline to accommodate variable dataset dimensions and explore scalable attention mechanisms for larger spatial domains [cao2022swin].
Statistical model.
We adopt a Gaussian process framework via SAR approximation, which enforces monotonic decay of the covariance and cannot capture global teleconnections or nonlinear dynamics such as eddies. Our current formulation omits explicit modeling of an additional white noise process, and could benefit from increased smoothing in areas with long-range correlation structures. Extensions to this work could explore multi-resolution structures [katzfuss2017multi, frkv2, latticekrig], estimate a spatially varying noise term, and approximate nonlocal dependence patterns that arise in physical systems.
While our framework enables efficient simulation of plausible ensembles, it ultimately inherits the assumptions and constraints of the underlying statistical model. Applications beyond climate model emulation—such as epidemiology, hydrology, or materials science—are feasible, although they may require tailoring the data generation strategy and retraining.

### v2

@@ 4 PARAMETER ESTIMATION NETWORKS
We adapt three I2I networks for parameter estimation: a fully convolutional U-Net, a modified ViT, and a hybrid network inspired by the TransUNet architecture. A range of CNNs representative of the local neural estimator literature serve as our baselines.
UNet
We use a standard U-Net [60] with a symmetric encoder-bottleneck-decoder structure and skip connections that propagate spatial information across resolutions. Our implementation replaces ReLU with GELU activations [24], employs group normalization [83], and omits dropout [71]. The number of channels in the bottleneck matches the transformer embedding dimension used in the transformer-based I2I networks.
ViT
The original ViT [16] was designed for image classification, so we remove classification-specific components such as the class token and final MLP head, and use a learnable linear layer to project the transformer output back to the original image resolution. The original 1D positional embeddings are replaced with a range of 2D alternatives; see Section 5.
STUN
Our hybrid architecture, a spatial TransUNet (STUN), is based upon a network originally designed for image segmentation [10]. It blends the two I2I networks discussed above, originally combining a pretrained convolutional encoder with a vision transformer and a shallow decoder. We retain the overall structure, but replace the encoder and decoder with the symmetric U-Net components used in our fully convolutional network.
CNN
Our local neural estimators follow the standard design of convolutional layers and a max pooling layer followed by an MLP. We use three networks with varying receptive window sizes that are representative of the local estimation literature [1, 34, 20, 64, 82]. The key difference is our networks have more parameters (1-2.5M), as compared to typical local estimators (50-700k).
@@ 4.1 Implementation and Training Details
All networks are trained with a batch size of $b=64$ and varying numbers of replicate input fields $M\in\{1,5,15,30\}$. Unless otherwise noted, all reported network sizes and metrics correspond to $M=30$ replicates. For the I2I networks, each training example consists of $M$ input fields of shape $[b,M,H,W]=[64,M,192,288]$, with corresponding output parameter fields $[64,3,192,288]$. For the local CNN estimators, we vary the receptive field $h=w=\{9,17,25\}$, with inputs of shape $[64,M,h,w]$ and scalar outputs $[64,3]$. We append the window size to the name of the CNN, thus $h=w=25$ is denoted CNN25. To encourage permutation invariance, we randomly shuffle the replicates across the channel dimension each training step.
Certain data augmentation techniques do not preserve the statistical structure of the fields. For example, image rotation does not preserve the angle of anisotropy $\theta$. Thus, we limit augmentation to spatial translation and random field negation. We generate datasets consisting of 8000 (I2I) and 80000 (CNN) samples with 30 replicates, and employ a 90/8/2 (train/validation/test) split, resulting in a test set of 160 samples for the I2I dataset, which both local and global methods are evaluated on. More details on implementation, storage, and software can be found in Appendix C.
@@ 5 SIMULATED DATA EXPERIMENTS
We generate a test set of $160$ samples following the procedure in Section 3. Each sample is a small ensemble $Y\in\mathbb{R}^{M\times 192\times 288}$ of $M$ replicates and an associated parameter image $\Phi\in\mathbb{R}^{3\times 192\times 288}$. For the I2I networks (UNet, ViT, STUN) we simply forward–propagate $Y$ to obtain $\hat{\Phi}_{\text{I2I}}$. For the local CNN baselines, we employ a pixel-by-pixel approach: we translate the CNN window across the field with a stride of 1, assigning the prediction to the central pixel of the window to build $\hat{\Phi}_{\text{CNN}}$. We use reflection padding—improving upon prior works which use zero padding—thus reducing edge-artifacts. Figure 4 contrasts an example $\Phi$ with estimates from the best performing global (STUN) and local (CNN25) networks. MLE is not attempted; local likelihood evaluation for one field would require hundreds of hours, exceeding the total runtime of our approach by orders of magnitude [82].
Positional Embeddings
The spatial nature of our problem leads us to study the effects of positional embeddings. We evaluate four embeddings for the transformer: none, 2-D sinusoidal [52], learned 2-D [16], and rotary (RoPE) [74]. RoPE yields the best performance across the majority of metrics for both ViT and STUN, and is adopted henceforth (see Appendix Table 2).
Number of Replicates
@@
Figure 4: True parameters $\Phi$ (left) are encoded into the LatticeKrig SAR model to simulate a testing sample $Y$, of which one replicate $\mathbf{y}^{(0)}$ is displayed (top). $Y$ is used as an input to STUN and a sliding window, local estimation strategy using CNN25, resulting in $\hat{\Phi}_{\text{STUN}}$ (middle), and $\hat{\Phi}_{\text{CNN25}}$ (right).
Table 1: Root mean square error (RMSE) for parameter estimation on a simulated test set using 1 and 30 replicates (reps). “Size” is the net parameter count (in millions). “Train” is the wall-clock time for training until early stopping, and “Eval” is the average time (out of 5) to process the $160$ sample test dataset. Arrows indicate desirable direction, and bold values indicate best performance.
[…]
@@
Our experiments highlight three consistent trends. (i) Global I2I networks pose a significant improvement over local CNNs: STUN and UNet often display 4-5x lower RMSE and are almost unaffected by shrinking the ensemble size, whereas CNNs are noticeably sensitive. (ii) Adding attention helps only marginally in this setting, and attention on its own lags behind approaches that include multiscale convolutions. STUN edges out UNet but at 4x the parameter count. ViT lags behind the other I2I networks, perhaps as it requires more training data [16]. (iii) Global I2I networks take longer to train due to a higher parameter count, yet are amortized at inference time. In order to achieve the results of a single forward pass through an I2I network, the local estimators must perform $H\times W=55,296$ forward passes. Consequently, I2I networks perform inference 100–1000 times faster than local neural estimators.
@@ 6 CLIMATE APPLICATION
We evaluate our framework by estimating parameters from climate model outputs, using these parameters to generate large, synthetic ensembles, and then comparing the quality of the I2I-based and local CNN-based emulators. We consider ensembles of surface temperature sensitivity fields from three climate models with differing numbers of replicates and resolutions: MPI-ESM (50 members, $192\times 96$ resolution, [51]), CESM1 (30, $288\times 192$, [28]), and EC-Earth3 (72, $512\times 256$, [17]). While diffusion networks have recently shown promise in adjacent emulation applications [61, 4], only having 30-72 fields per model renders the training of generative models impractical in this setting.
@@
The fields represent the local changes in temperature given an increase of $1^{\circ}$C in global temperature. Due to the chaotic nature of the ocean-atmosphere system, we can treat each field within an ensemble as an independent replicate sampled from the model’s “true” climate sensitivity. The data is preprocessed into standardized temperature sensitivity anomalies where each pixel has zero mean and unit variance. We then perform parameter estimation with all global I2I and local CNN estimators on 30 randomly selected fields from each ensemble. Regardless of the estimation method used, the LatticeKrig SAR model simulates $1000$ fields in less than one minute on a single laptop CPU, a stark contrast to the tens of millions of core hours required to generate the original ensembles.
@@
Figure 5: Top: Standardized temperature fields drawn from the CESM1 ensemble (left), the STUN-based emulator (middle), and the CNN25-based emulator (right). Bottom: Correlations with a chosen location in the Niño 3.4 region at ($212^{\circ}$E, $1^{\circ}$N) for the same three ensembles. The STUN-based emulator better preserves spatial relationships, including the zonal correlation structure along the equator and the meridional oceanic correlation range.
For each climate model, we perform the following experiment: We use the I2I-based and CNN-based emulators to generate 1000 fields each, and evaluate how well these synthetic ensembles preserve the spatial relationships present in the climate model outputs. Absolute prediction error is not meaningful here as the “truth” is itself a Monte-Carlo sample of internal climate variability. Thus, we compare the second-order structure through a covariance analysis. Rather than computing the entire correlation matrices of the true and simulated fields, which can get as large as $131,072\times 131,072$ (EC-Earth3), we compare a representative sample of $50$ rows.
First, we randomly select 50 anchor locations. We then calculate each anchor’s Pearson correlation with every other location for the original fields, and those from the I2I-based and CNN-based emulators. The average RMSE between the 50 true rows and their simulated counterparts is then calculated. We find that I2I-based emulators consistently outperform CNN-based emulators in capturing spatial relationships, exhibiting significantly lower RMSEs in eight of nine cases (paired t-tests; Appendix Table 5).
@@
As a representative qualitative comparison, we visualize empirical correlations in CESM1 with a chosen point in the Niño 3.4 region—a region of the tropical Pacific most correlated with the El Niño–Southern Oscillation phenomenon [3]—for the STUN-based and CNN25-based emulators (Figure 5). The STUN-based emulator better preserves the expected zonal east-west correlation structure along the equator with a more realistic correlation range. The CNN-based emulator systematically over-smooths and inflates the oceanic correlation range, particularly in the meridional north-south direction. The original ensemble correlation field is noisier as the CESM1 ensemble contains only 30 members; whereas our synthetic ensembles contain 1,000. These patterns remain consistent across climate models and network pairs. In sum, the I2I-based methods outperform the local CNN-based methods both quantitatively and qualitatively in representing the underlying correlation structure of climate sensitivity fields.
@@ 7 RELATED WORKS
Neural Parameter Estimation
In situations where the likelihood function is intractable, but simulation from the model is feasible, neural networks have emerged as a powerful alternative to MLE [38, 87]. Neural parameter estimation combines ideas from simulation-based inference [13, 69] and learning to optimize (L2O) [12, 84] by training networks to identify maxima of the Bayes risk [64]. The training cost is then amortized by repeated use. Until now, these approaches have been limited to local estimation [1, 34, 64, 56, 77, 20, 57], typically using multi-layer perceptrons (MLPs) and CNNs. This line of research is the main inspiration for this work, which advances the field with simultaneous, non-stationary parameter inference and the use of I2I networks.
Climate Model Applications
While climate model emulation is not the sole application of our method, it provides an illustrative example. Physics-based climate models require $10^{7}$$10^{8}$ core hours, necessitating data-driven emulation [18]. Despite recent advances in deterministic [54, 33, 47, 5, 45, 29, 11] and ensemble-based probabilistic [31, 55, 35, 67] weather forecasting, comparatively fewer methods focus explicitly on long-term climate projections [32, 26]. Our approach draws ideas from statistical methods, which explicitly link parameter estimation with emulation [7, 70, 49, 82, 8], and machine learning (ML) methods, both deterministic [46, 80, 9] and probabilistic [61, 62] which emulate directly from initial forcings or prior timesteps. Specifically, we combine the straightforward uncertainty quantification and interpretable parameters from the statistical model with a deterministic ML approach for efficient parameter estimation. Purely ML-based climate emulators typically require an extensive corpus of training data [79, 85] and can exhibit limited generalization capabilities beyond their training distributions [25]. We sidestep this by generating synthetic training data that mimics non-stationarity in geophysical settings. Our framework does not serve as a replacement for ESMs, but rather a complementary method for augmenting ensemble sizes, and reducing the number of “ground truth” ESM runs that must be computed.
@@ 8 LIMITATIONS AND EXTENSIONS
Estimation Networks
Our I2I networks assume complete, regularly gridded data. Replicate count is also fixed at training time, with only a soft constraint to enforce permutation invariance. A hard constraint via aggregation [86] might resolve this, and warrants further exploration. Transformer-based networks would likely benefit from larger datasets or advanced training techniques such as student-teacher distillation [76] and a larger dataset. Future work could extend the data generation pipeline to accommodate variable training dataset dimensions and explore scalable attention mechanisms for larger spatial domains [6]. Our network choices establish a baseline, but there exist many additional architectures for future work to explore [39, 58, 30, 78, 2].
Statistical Model
We adopt a Gaussian process framework via SAR approximation, which enforces monotonic decay of the covariance and cannot capture teleconnections or nonlinear dynamics such as eddies. Our current formulation omits explicit modeling of an additional white noise process, and could benefit from increased smoothing in areas with long-range correlation structures. Extensions to this work could explore multi-resolution structures [27, 63, 48], estimate a spatially varying noise term, and approximate nonlocal dependence patterns that arise in physical systems.
While our framework enables efficient parameter estimation and simulation of plausible ensembles, it ultimately inherits the assumptions and constraints of the underlying statistical model. Applications across a broad range of fields such as epidemiology, hydrology, or materials science are feasible, although they may require tailoring the data generation strategy and retraining.